A force of 23N is applied to the wire and it extends by 0.20mm. Find the Young modulus of the copper.
1
Expert's answer
2020-11-17T13:03:21-0500

"Y\frac{\Delta l}{l}=\frac{F}{A}=\frac{F}{0.25\pi d^2}\\Y\frac{0.0002}{1}=\frac{23}{0.25\pi (0.0011)^2}\\Y=1.2\cdot10^{11}\frac{N}{m^2}"